# Strategic Effort and Bandwagon Effects in Finite Multi-Stage Games with Non-Linear Externalities: Evidence from Triathlon using a Leave-One-Out IV Strategy

### v1

@@ A.3 Robustness: Leave-one-out Peer-IV 2SLS Estimates Using Theory-Implied Drafting Benefit Function
(1) Group Size $>$ 1 (2) Group Size $<$ 20 (3) Group Size $<$ 10 Drafting Benefit (Fit) -6.650∗∗∗ -10.662∗∗∗ -16.711∗∗ (0.654) (1.871) (5.239) Leader -2.845∗∗∗ -3.676∗∗∗ -3.980∗∗ (0.288) (0.659) (1.256) Observations 45,215 29,073 21,190 Athletes (FE) 15,762 12,544 10,534 Events (FE) 779 789 786 Clusters (FE) 149 149 149 RMSE 1.717 2.570 3.165 Adj. $R^{2}$2.100 -6.635 -12.100 Within $R^{2}$7.186 -22.200 -42.000 First Stage F-stat 254.7 53.0 16.3 First Stage $p$-value $<2.2\times 10^{-16}$ $3.54\times 10^{-13}$ $5.41\times 10^{-5}$ Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Each column reports two-stage least squares (2SLS) estimates of $\log(r_{i}+1)$, where $r_{i}$ is the athlete’s centered within-event rank. The endogenous variable is a theory-based drafting benefit: $B(d_{i})=\begin{cases}0,&\text{if }d_{i}\leq 3\\ \gamma\left(1-e^{-\lambda(d_{i}-3)}\right),&\text{if }d_{i}>3\end{cases}$ with $\gamma=1$ and $\lambda=0.5$. The instrument is the cluster-level peer mean swim time (leave-one-out). Fixed effects for athlete, cluster, and event are included. Standard errors are clustered at the event level. In Column (1), covering all valid group sizes, the drafting benefit significantly reduces disutility (better performance), with a very strong first-stage instrument. Effects are slightly larger in magnitude in Columns (2) and (3), reflecting increased sorting and clearer benefit signals in smaller, more cohesive drafting groups. The leader dummy is negative and significant in all samples, suggesting that leading is correlated with stronger athletes, consistent with selection or effort asymmetries. Log-rank values may include centering; NA values and log issues led to removal of 78,384; 56,239; and 42,541 observations in Columns (1) to (3) respectively. Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
@@
(1) (2) (3) (4) (5) (6) Group $>$ 1 Group > 1, No Leader Group $<$ 20 Group $<$ 20, No Leader Group $<$ 10 Group $<$ 10, No Leader Theoretic Drafting Benefit Fit $B(d_{i})$ $-6.650^{***}$ $-2.871^{***}$ $-4.836^{***}$ $-3.609^{***}$ $-9.615^{***}$ $-7.069^{***}$ (0.654) (0.267) (0.697) (0.468) (2.371) (1.509) Leader Indicator $-2.845^{***}$ $-1.645^{***}$ $-2.053^{***}$ (0.288) (0.242) (0.511) Observations 45,215 34,061 22,026 22,026 16,035 16,035 Athletes (FE) 15,762 12,181 10,075 10,075 8,408 8,408 Events (FE) 779 438 498 498 523 523 Clusters (FE) 149 178 178 178 178 178 RMSE 1.717 0.925 1.201 1.020 1.661 1.339 Adj. $R^{2}$2.10 0.483 -0.075 0.225 -1.383 -0.550 Within $R^{2}$7.19 -1.75 -5.13 -3.42 -13.20 -8.21 First-stage $F$-stat 254.7 513.7 99.1 148.5 23.4 37.4 Wu–Hausman stat 1,901.0 918.2 509.1 508.6 298.5 297.7 Athlete FE Event FE Cluster FE Notes: 2SLS estimates of $\log(r_{i}+1)$, where $r_{i}$ is centered rank. The endogenous regressor is the theory-based drafting benefit: $B(d_{i})=1-e^{-\lambda(d_{i}-3)}$ for $d_{i}>3$, and 0 otherwise, with $\lambda=0.5$. Instrument: leave-one-out peer swim time. Fixed effects for athlete, event, and cluster are included. Standard errors clustered at the event level. Leader coefficients are shown where included. Column labels correspond to different subsamples and whether the leader dummy is included. Drafting benefits are consistently significant and negative, with stronger effects in smaller groups. Significance levels: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$
Appendix B Data and Descriptive Statistics as in (Reichel , 2025)
@@ Appendix D Pooled Bandwagon Effects of Latter Drafting Positions on Overall Rank
Because $\text{Treat}_{it}$ may be endogenous—potentially correlated with $\varepsilon_{it}$—we instrument for it using the exogenous variation in peer group swim ability. The first-stage equation is:
$\text{Treat}_{it}=\pi\cdot\text{PeerSwimMean}_{it}+\eta_{i}+\theta_{t}+u_{it}$ (2)
@@
  • $\text{PeerSwimMean}_{it}$ is the leave-one-out average swim score of athlete $i$’s drafting cluster in event $t$.
@@
Under the standard IV assumptions—instrument relevance ($\pi\neq 0$), instrument exogeneity ($\mathbb{E}[\text{PeerSwimMean}_{it}\cdot\varepsilon_{it}]=0$), and monotonicity—the parameter $\beta$ captures the local average treatment effect (IV-LATE) of being a drafter in latter pooled drafter position band on log-transformed, adjusted rank.

### v2

@@ A.3 Robustness: Leave-one-out Group-IV 2SLS Estimates Using Theory-Implied Drafting Benefit Function
(1) Group Size $>$ 1 (2) Group Size $<$ 20 (3) Group Size $<$ 10 Drafting Benefit (Fit) -6.650∗∗∗ -10.662∗∗∗ -16.711∗∗ (0.654) (1.871) (5.239) Leader -2.845∗∗∗ -3.676∗∗∗ -3.980∗∗ (0.288) (0.659) (1.256) Observations 45,215 29,073 21,190 Athletes (FE) 15,762 12,544 10,534 Events (FE) 779 789 786 Clusters (FE) 149 149 149 RMSE 1.717 2.570 3.165 Adj. $R^{2}$2.100 -6.635 -12.100 Within $R^{2}$7.186 -22.200 -42.000 First Stage F-stat 254.7 53.0 16.3 First Stage $p$-value $<2.2\times 10^{-16}$ $3.54\times 10^{-13}$ $5.41\times 10^{-5}$ Wu-Hausman $p$-value $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ $<2.2\times 10^{-16}$ Notes: Each column reports two-stage least squares (2SLS) estimates of $\log(r_{i}+1)$, where $r_{i}$ is the athlete’s centered within-event rank. The endogenous variable is a theory-based drafting benefit: $B(d_{i})=\begin{cases}0,&\text{if }d_{i}\leq 3\\ \gamma\left(1-e^{-\lambda(d_{i}-3)}\right),&\text{if }d_{i}>3\end{cases}$ with $\gamma=1$ and $\lambda=0.5$. The instrument is the cluster-level Group mean swim time (leave-one-out). Fixed effects for athlete, cluster, and event are included. Standard errors are clustered at the event level. In Column (1), covering all valid group sizes, the drafting benefit significantly reduces disutility (better performance), with a very strong first-stage instrument. Effects are slightly larger in magnitude in Columns (2) and (3), reflecting increased sorting and clearer benefit signals in smaller, more cohesive drafting groups. The leader dummy is negative and significant in all samples, suggesting that leading is correlated with stronger athletes, consistent with selection or effort asymmetries. Log-rank values may include centering; NA values and log issues led to removal of 78,384; 56,239; and 42,541 observations in Columns (1) to (3) respectively. Significance levels: ${}^{***}p<0.01$, ${}^{**}p<0.05$, ${}^{*}p<0.1$
@@
(1) (2) (3) (4) (5) (6) Group $>$ 1 Group > 1, No Leader Group $<$ 20 Group $<$ 20, No Leader Group $<$ 10 Group $<$ 10, No Leader Theoretic Drafting Benefit Fit $B(d_{i})$ $-6.650^{***}$ $-2.871^{***}$ $-4.836^{***}$ $-3.609^{***}$ $-9.615^{***}$ $-7.069^{***}$ (0.654) (0.267) (0.697) (0.468) (2.371) (1.509) Leader Indicator $-2.845^{***}$ $-1.645^{***}$ $-2.053^{***}$ (0.288) (0.242) (0.511) Observations 45,215 34,061 22,026 22,026 16,035 16,035 Athletes (FE) 15,762 12,181 10,075 10,075 8,408 8,408 Events (FE) 779 438 498 498 523 523 Clusters (FE) 149 178 178 178 178 178 RMSE 1.717 0.925 1.201 1.020 1.661 1.339 Adj. $R^{2}$2.10 0.483 -0.075 0.225 -1.383 -0.550 Within $R^{2}$7.19 -1.75 -5.13 -3.42 -13.20 -8.21 First-stage $F$-stat 254.7 513.7 99.1 148.5 23.4 37.4 Wu–Hausman stat 1,901.0 918.2 509.1 508.6 298.5 297.7 Athlete FE Event FE Cluster FE Notes: 2SLS estimates of $\log(r_{i}+1)$, where $r_{i}$ is centered rank. The endogenous regressor is the theory-based drafting benefit: $B(d_{i})=1-e^{-\lambda(d_{i}-3)}$ for $d_{i}>3$, and 0 otherwise, with $\lambda=0.5$. Instrument: leave-one-out Group swim time. Fixed effects for athlete, event, and cluster are included. Standard errors clustered at the event level. Leader coefficients are shown where included. Column labels correspond to different subsamples and whether the leader dummy is included. Drafting benefits are consistently significant and negative, with stronger effects in smaller groups. Significance levels: ${}^{***}p<0.001$, ${}^{**}p<0.01$, ${}^{*}p<0.05$
Appendix B Data and Descriptive Statistics (as in (Reichel , 2025))
@@ Appendix D Pooled Bandwagon Effects of Latter Drafting Positions on Overall Rank
Because $\text{Treat}_{it}$ may be endogenous—potentially correlated with $\varepsilon_{it}$—we instrument for it using the exogenous variation in Group group swim ability. The first-stage equation is:
$\text{Treat}_{it}=\pi\cdot\text{GroupSwimMean}_{it}+\eta_{i}+\theta_{t}+u_{it}$ (2)
@@
  • $\text{GroupSwimMean}_{it}$ is the leave-one-out average swim score of athlete $i$’s drafting cluster in event $t$.
@@
Under the standard IV assumptions—instrument relevance ($\pi\neq 0$), instrument exogeneity ($\mathbb{E}[\text{GroupSwimMean}_{it}\cdot\varepsilon_{it}]=0$), and monotonicity—the parameter $\beta$ captures the local average treatment effect (IV-LATE) of being a drafter in latter pooled drafter position band on log-transformed, adjusted rank.